A force of 6N acts horizontally on a stationary mass of 3kg for 6 seconds.the kinetic energy gainesd by the mass in joules is
1
Expert's answer
2021-09-28T11:12:13-0400
"a=\frac{6}{3}=2\frac{m}{s^2}\\s=0.5(2)(6)^2=36\ m\\E=6(36)=216\ J"